# Hal-Eval: A Universal and Fine-grained Hallucination Evaluation Framework for Large Vision Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) exhibit remarkable capabilities but struggle with "hallucinations"—inconsistencies between images and their descriptions. Previous hallucination evaluation studies on LVLMs have identified hallucinations in terms of objects, attributes, and relations but overlooked complex hallucinations that create an entire narrative around a fictional entity. In this paper, we introduce a refined taxonomy of hallucinations, featuring a new category: Event Hallucination. We then utilize advanced LLMs to generate and filter fine-grained hallucinatory data consisting of various types of hallucinations, with a particular focus on event hallucinations, laying the groundwork for integrating discriminative and generative evaluation methods within our universal evaluation framework. The proposed benchmark distinctively assesses LVLMs’ ability to tackle a broad spectrum of hallucinations, making it a reliable and comprehensive tool for gauging LVLMs’ efficacy in handling hallucinations. We will release our code and data.
Hal-Eval: A Universal and Fine-grained Hallucination Evaluation Framework for Large Vision Language Models
[…]
1 Introduction
@@
Large Language Models (LLMs) such as GPT-4 OpenAI (2023), LLaMA Touvron et al. (2023a), and LLaMA2 Touvron et al. (2023b) have markedly enhanced capabilities in natural language understanding (NLU) and generation (NLG). Building on these advancements, recent Large Vision-Language Models (LVLMs) have shown increased proficiency in handling both textual and visual information, sparking significant interest among researchers. Zhu et al. (2023); Liu et al. (2023e, c); Anil et al. (2023); Dai et al. (2023); Ye et al. (2023); Chen et al. (2023a); Li et al. (2023d).
Figure 1: Different types of hallucination. Event hallucination, which involves more complex vision-language discrepancy compared to other types of hallucination, is commonly overlooked by previous efforts.
@@
Despite the promising developments in LVLMs, they broadly face the pivotal obstacle of hallucination, which refers to the discrepancy between factual content in images and the associated generated textual descriptions. As hallucination poses significant concerns for the LVLMs’ reliability and robustness Liu et al. (2024); Li et al. (2023e); Liu et al. (2023a); Sun et al. (2023a); Guan et al. (2023); Zhai et al. (2023), researchers have devised strategies for hallucinations evaluation to bolster the practical deployment of LVLMs, including discriminative and generative methods. The former method directly prompts candidate LVLMs to determine the presence of a particular hallucination, whereas the latter assesses the text produced by these candidate LVLMs.
@@
Prior research Liu et al. (2024); Jing et al. (2023); Gunjal et al. (2023) has delineated vision-language mismatches as issues of non-existent objects, incorrect object attributes, or inaccurate object relations, yet it does not encompass the full spectrum of hallucinations observed in LVLMs. For example, as depicted in Figure 1, LVLM outputs can exhibit more intricate hallucinations, such as "At the same time, a kangaroo is hopping near the zebras, seemingly foraging." This type of hallucination invents a fictional target and weaves an entire narrative around it, including its attributes, relationships, and actions. We categorize these intricate narratives as event hallucinations. Presently, there is an absence of fine-grained hallucination evaluation benchmarks for LVLMs that comprehensively address the various types of hallucinations—such as artificial objects, relationships, attributes, and events—while also accommodating both discriminative and generative evaluation methods.
@@
To this end, we first devise a strategy for creating datasets with fine-grained hallucination instances. By leveraging the advanced capabilities of LLMs for generating and filtering hallucinatory data, we develop an efficient and cost-effective pipeline to assemble samples featuring hallucinations of objects, relationships, attributes, and events. This pipeline then serves as a solid foundation for unifying discriminative and generative evaluation methodologies in our framework:
@@
  • For the generative evaluation, our pipeline facilitates the creation of a large-scale hallucinatory dataset. This dataset serves to fine-tune an LVLM into a specialized evaluator, namely Hal-Evaluator. This evaluator assesses LVLM-generated descriptions and their associated images, identifying various hallucination types without needing reference captions.
We conduct thorough experiments and analysis with five leading LLMs across in- and out-of-domain datasets within our framework, assessing their performance in terms of hullucination under both discrimitive and generative paradigm. Our key findings are:
  • Our method mitigates the observed bias in evaluated LVLMs, which often default to "Yes" responses to discriminative questions in benchmarks like POPE, promoting more balanced responses.
@@
  • The hallucinatory samples used to train our evaluator also serve as effective supervised fine-tuning data for LVLMs, contributing to the reduction of hallucinations and enhancement of their benchmark performance.
@@
In summary, we introduce a novel hallucination category (event hallucination) of LVLMs, a fine-grained evaluation framework for LVLMs that spans various hallucination types and unify discriminative and generative approaches, along with some groundbreaking insights to guide future research on vision-language hallucination.
Figure 2: The sub-figure (a) illustrates the pipeline of AFHA. The lower left sub-figure (b) visualizes the prompt used for generating Object Hallucination Annotations (Please refer to Appendix B.2 for the prompt of other hallucination annotations.), while the lower right sub-figure visualizes examples of annotations for four types of hallucinations.
2 AFHA: Automatic Fine-grained Hallucination Annotation Pipline
The existing multimodal hallucination research suffers from a lack of large-scale datasets with fine-grained annotations specific to hallucinations. To address this issue, we design AFHA, an Automatic Fine-grained Hallucination Annotation pipline featuring annotations for four hallucination types and specific hallucination content.
Data Annotation. We annotated image-text paired data using the GPT-4 prompt method. We initially established a rigorous definition for various types of hallucinations. Building upon this groundwork, we engaged GPT-4 to rephrase the collated image-text pairs in line with the diverse classifications of hallucinations. This step involved inject distinctive hallucinatory elements into the original captions. An example of a prompt designed to generate ’object’ annotations is illustrated in Figure 2(b). The outcome of this procedure was a collection of image descriptions enriched with specified hallucination categories. Moreover, we delegated to GPT-4 the responsibility of concocting specific hallucinatory content. Consequently, this strategy yielded an extensively annotated dataset facilitated by GPT-4, with samples of annotations across different types of hallucinations displayed in Figure 2(c). For more details, please refer to the Appendix B.2.
Data Filtering. Following the initial annotation phase, we identified that the quality of the labeled data remained unsatisfactory. Random sampling revealed that approximately 30% of the annotated dataset still harbored noise that failed to meet our stringent labeling criteria. Hence, we proceeded to craft a tailored prompt to commission GPT-4 for the task of purging and refining the noisy annotations, a process thoroughly outlined in the Appendix B.3. Subsequent to GPT-4’s meticulous cleanup operation, a manual verification process ascertained that over 97% of the data accorded with the stipulated annotation standards.
3 Hal-Eval: A Universal Hallucination Evaluation Benchmark
Building upon AFHA, we have proposed a comprehensive and universal hallucination evaluation benchmark named Hal-Evaluator. As shown in Figure 3, Hal-Evaluator includes both Discriminative Evaluation and Generative Evaluation, and is capable of effectively evaluating different types of hallucinations. In the below subsections, we will first introduce the evaluation datasets used in discriminative evaluation, then we will introduce detailed descriptions of discriminative evaluation. Finally, we will introduce generative evaluation.
3.1 Discrimniative Evaluation
3.1.1 Constructing Evaluation Dataset
Data Collection: In the past, evaluation benchmarks such as POPE Li et al. (2023e) have mainly relied on datasets from manually annotated sources like COCO Lin et al. (2014). However, these datasets are extensively used in the training and fine-tuning of many LVLMs, resulting in the evaluation data being in-domain with these models’ training data. Consequently, the models’ zero-shot hallucination capabilities cannot be genuinely assessed. To address this, our evaluation dataset is split into two parts. One part consists of in-domain evaluation data, composed of image-text pairs from COCO 2014 validation and COCO 2017 test sets. The other part is sourced from web-based datasets such as CC Changpinyo et al. (2021), SBU Ordonez et al. (2011), and LAION Schuhmann et al. (2022), providing out-of-domain data sampled randomly.
Data Annotation: We leveraged the automatic annotation process AFHA scripted in Section 2 to annotate both in-domain and out-of-domain evaluation data. This process resulted in 5,000 detailed annotations for in-domain evaluation data and 5,000 for out-of-domain evaluation data, each delineating hallucination types and content, as visually depicted in the sample in Sub-Figure 2 (c). A formalized annotated sample is represented as $S=\{I,C^{T},C^{O},C^{R},C^{E},C^{A}\}$ where I denotes the image, $C^{T}$ represents the correct image caption, and $C^{O},C^{R},C^{E},C^{A}$ respectively denote captions containing Object hallucinations, Relation hallucinations, Event hallucinations, and Attribute hallucinations.
3.1.2 Evaluation Process
In prior work, the discriminative evaluation method proposed for evaluating a certain type of hallucination directly asked LVLMs if the content of that type existed in the image. For instance, evaluating object hallucinations directly inquired about the presence of a specific object. In contrast, we have proposed a more natural questioning method as following.
@@ 3.1.2 Evaluation Process
Figure 3: This figure provides a schematic of the discriminative evaluation and generative evaluation used in Hal-Eval.
@@
Evaluation Metric. Similar to POPE Li et al. (2023e), we also use Accuracy, Precision, Recall, F1 score, and "Yes" ratio as the evaluation metrics. Here, Accuracy represents the number of correctly answered cases, while Precision and Recall respectively indicate the proportion of correctly answered questions with responses "Yes" or "No." F1 score integrates the outcomes of Precision and Recall, which we select as the primary evaluation metric. The "Yes ratio" serves as a reference for analyzing model behaviors.
3.2 Generative Evaluation
3.2.1 Overview
@@
Regarding generative evaluation, current evaluation methods either rely on proprietary models that require subscription fees, such as GPT-4, or depend on fine-tuned large language models (LLMs) that necessitate additional ground truth annotations—a process that is prohibitively expensive. This significantly restricts the scalability of evaluating models. In response, we propose Hal-Evaluator, a reference-free, open-source evaluation model designed specifically to detect hallucinatory content. Hal-Evaluator is fine-tuned on LLaVA 1.5 Liu et al. (2023c) which is also a LVLM; as illustrated in Figure 3, it takes as input the description of an image provided by the LVLMs under evaluation, as well as the corresponding image itself. Following this, it evaluate whether the description contains hallucinations. If hallucinations are detected, it provides the specific content and categorization of the hallucinations. Ultimately, it can even modify the hallucinated information in the description to output an accurate depiction. In this way, our generative evaluation eliminates the need for additional reference annotation, enabling hallucination evaluation based solely on the content of the image.
@@
To train the Hal-Evaluator which capable of effectively identifying different types of hallucinations, a large-scale, fine-grained hallucinatory image-text dataset is necessary as they facilitate the refinement of training for Hal-Evaluator intended to detect and correct hallucinatory content. However, there currently exists no dataset of this scale with detailed annotations. Therefore, we initially constructed Hal-Data, the first large-scale, fine-grained dataset with hallucination annotations, based on the AFHA pipeline.
3.2.2 Construction of Hal-Data

### v2

@@ Abstract.
Large Vision-Language Models (LVLMs) exhibit remarkable capabilities but struggle with ”hallucinations”—inconsistencies between images and their descriptions. Previous hallucination evaluation studies on LVLMs have identified hallucinations in terms of objects, attributes, and relations but overlooked complex hallucinations that create an entire narrative around a fictional entity. In this paper, we introduce a refined taxonomy of hallucinations, featuring a new category: Event Hallucination. We then utilize advanced LLMs to generate and filter fine-grained hallucinatory data consisting of various types of hallucinations, with a particular focus on event hallucinations, laying the groundwork for integrating discriminative and generative evaluation methods within our universal evaluation framework. The proposed benchmark distinctively assesses LVLMs’ ability to tackle a broad spectrum of hallucinations, making it a reliable and comprehensive tool for gauging LVLMs’ efficacy in handling hallucinations. We will release our code and data.
Hallucination, Large Vision Language Model, Event, Fine-grain
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680576isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
@@
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023a), LLaMA (Touvron et al., 2023a), and LLaMA2 (Touvron et al., 2023b) have markedly enhanced capabilities in natural language understanding (NLU) and generation (NLG). Building on these advancements, recent Large Vision-Language Models (LVLMs) have shown increased proficiency in handling both textual and visual information, sparking significant interest among researchers. (Zhu et al., 2023a; Liu et al., 2023e, c; Anil et al., 2023; Dai et al., 2023; Ye et al., 2023; Chen et al., 2023b; Li et al., 2023c).
@@
Despite the promising developments in LVLMs, they broadly face the pivotal obstacle of hallucination, which refers to the discrepancy between factual content in images and the associated generated textual descriptions. As hallucination poses significant concerns for the LVLMs’ reliability and robustness (Liu et al., 2024; Li et al., 2023a; Liu et al., 2023f; Wang et al., 2023b; Sun et al., 2023b; Guan et al., 2023; Zhai et al., 2023), researchers have devised strategies for hallucinations evaluation to bolster the practical deployment of LVLMs, including discriminative and generative methods. The former method directly prompts candidate LVLMs to determine the presence of a particular hallucination, whereas the latter assesses the text produced by these candidate LVLMs.
Figure 1. Different types of hallucination. Event hallucination, which involves more complex vision-language discrepancy compared to other types of hallucination, is commonly overlooked by previous efforts.
@@
Prior research (Liu et al., 2024; Jing et al., 2023; Gunjal et al., 2023) has delineated vision-language mismatches as issues of non-existent objects, incorrect object attributes, or inaccurate object relations, yet it does not encompass the full spectrum of hallucinations observed in LVLMs. For example, as depicted in Figure 1, LVLM outputs can exhibit more intricate hallucinations, such as ”At the same time, a kangaroo is hopping near the zebras, seemingly foraging.” This type of hallucination invents a fictional target and weaves an entire narrative around it, including its attributes, relationships, and actions. We categorize these intricate narratives as event hallucinations. Our preliminary experiments indicate that the occurrence of event hallucinations significantly escalates as the output length of LVLMs increases, underscoring its significance as a hallucination phenomenon that warrants attention. However, there is an absence of fine-grained hallucination evaluation benchmarks for LVLMs that comprehensively address the various types of hallucinations—such as artificial objects, relationships, attributes, and events—while also accommodating both discriminative and generative evaluation methods.
@@
To this end, we propose a universal, fine-grained hallucination evaluation framework for LVLMs. This framework comprehensively evaluates a broad spectrum of hallucination types, encompassing objects, relationships, attributes, and notably, events with discriminative and generative evaluation methodologies. Specifically, we first develop an automatic annotation pipeline for fine-grained hallucinations, which leverages the sophisticated capabilities of GPT4 to generate and filter hallucinatory data. This pipeline then serves as a solid foundation for unifying discriminative and generative evaluation methodologies in our framework:
@@
  • For the generative evaluation, our pipeline facilitates the creation of a large-scale hallucinatory dataset. This dataset serves to fine-tune an LVLM into a specialized evaluator, Hal-Evaluator. This evaluator assesses LVLM-generated descriptions and associated images, identifying various hallucination types without needing reference captions.
We conduct thorough experiments and analysis with six leading LLMs within our framework, assessing their performance in terms of hallucination under both discriminative and generative paradigms. Our key findings are:
  • The existing three categories of hallucinations (object, attribute, relation) overlook the existence of event-type hallucinations and are, therefore, insufficient to encompass all types of hallucinations.
@@
  • The suitability of evaluation methodology varies according to the type of hallucinations. Using discriminative and generative evaluations together gives a fuller view of tendencies to LVLM hallucination.
@@
  • The hallucinatory samples used to train our evaluator also serve as effective supervised fine-tuning data for LVLMs, contributing to reducing hallucinations and enhancing their benchmark performance.
@@
In summary, we introduce a novel hallucination category (event hallucination) of LVLMs, a universal and fine-grained evaluation framework for LVLMs that spans various hallucination types and unifies discriminative and generative approaches, along with some groundbreaking insights to guide future research on vision-language hallucination.
2. Related work
Large Vision Language Model: Based on LLMs, there are three principal approaches to constructing LVLMs, all demonstrating potential for robust zero-shot generalization in the vision-language field. For instance, Flamingo (Alayrac et al., 2022) utilizes a fixed vision encoder paired with a sizable language model featuring gated cross-attention mechanisms for cross-modality matching. Meanwhile, PaLM-E (Driess et al., 2023) incorporates visual features via linear layers directly into the pre-trained PaLM (Chowdhery et al., 2022) framework, which delivers strong performance across a spectrum of practical applications. This integration strategy is widely employed by models like LLaVA (Liu et al., 2023d), Shikra (Chen et al., 2023b), and others. However, generating long visual sequences remains a notable constraint of this technique. To mitigate this, BLIP-2 (Li et al., 2023b), inspired by DETR (Carion et al., 2020), conceived a Q-former that effectively condenses the length of visual feature sequences. This concept has since been reflected in Kosmos-1 (Huang et al., 2023), mPLUG-Owl (Ye et al., 2023), and MiniGPT-4 (Zhu et al., 2023b).
Hallucination Evaluation: LVLM hallucination benchmarks specifically aim at non-hallucinatory generation or hallucination discernment. These benchmarks are classified according to the type of evaluation approach they follow: Discriminative (Dis) or Generative (Gen). Discriminative Benchmark: POPE (Li et al., 2023a), NOPE (Lovenia et al., 2023), and CIEM (Hu et al., 2023) are examples of discriminative benchmarks. Each of these benchmarks exclusively directs attention towards object hallucinations and utilizes accuracy as their primary evaluation metric. The metric is calculated by querying the presence of objects within images and comparing the model’s responses with the ground truth. Generative Benchmarks: modern research predominantly accentuates generative benchmarks over discriminative ones. Whilst discriminative benchmarks focus mainly on object-level hallucinations, generative benchmarks widen their scope to encompass a more extensive range of hallucinations, such as attribute and relation hallucinations (Gunjal et al., 2023; Liu et al., 2023f; Jing et al., 2023; Wang et al., 2023c; Sun et al., 2023a). AMBER (Wang et al., 2023a) emerges as a holistic benchmark that concludes both generative and discriminative tasks.
3. Preliminary: Hallucination in LVLMs
Figure 2. The left sub-figure shows the ratios of various hallucinations in mPLUG-owl’s image descriptions with token lengths under 20. The right sub-figure presents these ratios for descriptions exceeding 20 tokens.
@@ 3. Preliminary: Hallucination in LVLMs
Figure 3. This figure provides a schematic of the discriminative evaluation and generative evaluation used in Hal-Eval.
Previous works have characterized misalignment of hallucination as claims of non-existent objects, incorrect object attributes, or inaccurate object relations. However, we find that this only partially covers the spectrum of hallucinations present in LVLMs. For instance, as shown in the left part of Figure 7, the outputs of LVLMs include more complex hallucinations: ” to chase after a squirrel that has invaded its playtime .” We refer to these complex hallucinations as event hallucinations. To further clarify the concept of different hallucinations, we provide strict definitions for four different types of hallucinations in this paper:
Object hallucination: The LVLM inaccurately describes an object that does not exist in the image. This could be a misidentification, where the model correctly detects an object’s presence but incorrectly labels it, or an additional non-existent object, where the model asserts the presence of an object that does not exist.
Attribute hallucination: The LVLM correctly identifies an object that exists in the image, but inaccurately describes the attributes of that object. Attributes could include color, size, shape, position, or any other characteristic that defines the object.
Relation hallucination: The LVLM incorrectly describes the relationship between two or more objects that clearly exists in the image. This could involve misrepresenting spatial relationships (e.g., describing an object as being on top of another when it’s actually beside it), functional relationships (e.g., stating that a person is riding a bicycle when they are standing next to it), or other types of interactions or connections between objects.
Event hallucination: The LVLM not only describes a non-existent target but also constructs complete events around the non-existent target, including its attributes, relations, and actions. This type of hallucination involves a complex interplay of objects, attributes, and relations and often forms a narrative or sequence of actions that does not align with the actual content of the image.
Building upon these definitions, we further investigate the proportion of different types of hallucinations present within the output of LVLMs. As depicted in Figure 2, we collected 5,000 image-caption pairs from COCO (Lin et al., 2014). We had them described by mPLUG-owl (Ye et al., 2023) and LLaVA (Liu et al., 2023e), respectively. Subsequently, we provided both the ground truth image descriptions and the model-generated descriptions to GPT-4 (OpenAI, 2023a), prompting it to inspect whether these descriptions encompassed hallucinations and to categorize them based on Object, Attribute, Relationship, and Event hallucinations (For more details about the experiment, please refer to our Supplemental Material D.1.). We tabulated the proportions of different types of hallucinations at varying description lengths. As Figure 2 demonstrates, we noted a significant increase in the share of event hallucinations as the length of the description extends. This experimental observation substantiates our finding: The existing three categories of hallucinations (object, attribute, relation) overlook the existence of event-type hallucinations. They are, therefore, insufficient to encompass all types of hallucinations.
4. Method
We proposed a comprehensive and universal hallucination evaluation benchmark, Hal-Eval. As shown in Figure 3, Hal-Eval includes both Discriminative Evaluation and Generative Evaluation and can effectively evaluate different types of hallucinations. In the following subsection, we will initially introduce a fine-grained hallucination annotation pipeline. This pipeline is employed to construct both the evaluation dataset for Hal-Eval and a large-scale hallucination detection dataset known as Hal-Data. Hal-Data, in turn, serves as the training data for the Hal-evaluator, an evaluation model designed to perform generative hallucination evaluations. Subsequently, we will delve into an in-depth discussion of the discriminative evaluation process. Finally, we will elucidate the training procedure for the Hal-evaluator and outline how generative evaluations are conducted using this model.
4.1. Automatic Fine-grained Hallucination Annotation Pipeline
The existing multimodal hallucination research lacks large-scale datasets with fine-grained annotations specific to hallucinations. To address this issue, we design an automatic fine-grained hallucination annotation pipeline featuring annotations for four hallucination types and specific hallucination content.
Data Annotation. We annotated image-text paired data based on GPT-4. We initially established a rigorous definition for various types of hallucinations as we mentioned in Section 3. Building upon this groundwork, we engaged GPT-4 to rephrase the collated image-text pairs in line with the diverse classifications of hallucinations. This step involved injecting distinctive hallucinatory elements into the original captions. The outcome of this procedure was a collection of image descriptions enriched with specified hallucination categories. Moreover, we delegated to GPT-4 the responsibility of annotating the position of specific hallucinatory content in the image description. Please refer to the supplemental material C.2 for more details.
Data Filtering. Following the initial annotation phase, we identified that the quality of the labeled data remained unsatisfactory. Random sampling revealed that approximately 30% of the annotated dataset still harbored noise that failed to meet our stringent labeling criteria. Hence, we proceeded to craft a tailored prompt to commission GPT-4 for the task of purging and refining the noisy annotations, a process thoroughly outlined in the supplemental material C.3. Subsequent to GPT-4’s meticulous cleanup operation, a manual verification process ascertained that over 97% of the data accorded with the stipulated annotation standards.
4.2. Discriminative Evaluation
4.2.1. Constructing Evaluation Dataset
Data Collection: Previous benchmarks such as POPE (Li et al., 2023a) predominantly utilized manually annotated datasets like COCO (Lin et al., 2014). However, the COCO dataset is frequently employed to construct general benchmarks such as VQA v2 and visual grounding. These benchmarks are often used for instruction finetuning of LVLM models, which results in evaluation data being in the same domain as the models’ finetuning data. This overlap hinders a true assessment of the models’ zero-shot hallucination capabilities. To address this issue, we divided our evaluation dataset into two parts: in-domain data from the COCO 2014 validation and COCO 2017 test sets and an out-of-domain dataset randomly sampled from web-based data like CC (Changpinyo et al., 2021), SBU (Ordonez et al., 2011), LAION (Schuhmann et al., 2022).
Data Annotation: We used the automatic annotation process detailed in Section A to annotate both in-domain and out-of-domain evaluation datasets, resulting in 5,000 detailed annotations each. These annotations identify hallucination types and content. An annotated sample is represented as $S=\{I,C^{T},C^{O},C^{R},C^{E},C^{A}\}$, where I is the image, $C^{T}$ is the correct image caption, and $C^{O},C^{R},C^{E},C^{A}$ denote captions with Object, Relation, Event, and Attribute hallucinations, respectively.
4.2.2. Evaluation Process
In prior work, the discriminative evaluation method proposed for evaluating a specific type of hallucination asked LVLMs if the content of that type existed in the image. For instance, evaluating object hallucinations inquires about the presence of a specific object. In contrast, we have proposed a more natural questioning method, which is as follows:
@@ 4.2.2. Evaluation Process
Evaluation Metric. Similar to POPE (Li et al., 2023a), we also use Accuracy, Precision, Recall, F1 score, and ”Yes” ratio as the evaluation metrics. Here, Accuracy represents the number of correctly answered cases, while Precision and Recall, respectively, indicate the proportion of correctly answered questions with responses ”Yes” or ”No.” The F1 score integrates the outcomes of Precision and Recall, which we select as the primary evaluation metric. The ”Yes ratio” serves as a reference for analyzing model behaviors.
4.3. Generative Evaluation
4.3.1. Overview
@@
Regarding generative evaluation, current evaluation methods either rely on proprietary models that require subscription fees, such as GPT-4, or depend on fine-tuned large language models (LLMs) that necessitate additional ground truth annotations—a process that is prohibitively expensive. This significantly restricts the scalability of evaluating models. In response, we propose Hal-Evaluator, a reference-free, open-source evaluation model designed specifically to detect hallucinatory content. Hal-Evaluator is fine-tuned on LLaVA 1.5 (Liu et al., 2023c), which is also an LVLM; as illustrated in Figure 3, it takes as input the description of an image provided by the LVLMs under evaluation, as well as the corresponding image itself. Following this, it evaluates whether the description contains hallucinations. If hallucinations are detected, they provide the specific content and categorization of the hallucinations. Ultimately, it can even modify the hallucinated information in the description to output an accurate depiction. In this way, our generative evaluation eliminates the need for additional reference annotation, enabling hallucination evaluation based solely on the content of the image.
@@
To train the Hal-Evaluator, which is capable of effectively identifying different types of hallucinations, a large-scale, fine-grained hallucinatory image-text dataset is necessary as it facilitates the refinement of training for Hal-Evaluator intended to detect and correct hallucinatory content. However, no dataset of this scale with detailed annotations currently exists. Therefore, we initially constructed Hal-Data, the first large-scale, fine-grained dataset with hallucination annotations, based on the pipeline mentioned in Subsection A.
4.3.2. Instruction finetuning of Hal-Evaluator